OpenStudy (anonymous):

|-1 2 -3| |2 0 1| |3 -4 4|

OpenStudy (amistre64):

what to do with the matrix?

OpenStudy (anonymous):

evaluate the determinant

OpenStudy (amistre64):

ive got a nice shortcut to a 3x3 that is similar to a 2x2

OpenStudy (amistre64):

copy the first 2 columns on the end of it, and multiply the diags up and down; top goes negative and bottom goes positive, then add the results

OpenStudy (anonymous):

\[\left[\begin{matrix}-1 & 2 & -3 \\ 2 & 0 & 1 \\ 3 &-4 &4\end{matrix}\right]=-1\left[\begin{matrix}0 & 1 \\ -4 & 4\end{matrix}\right]-2\left[\begin{matrix}2 & 1 \\ 3 & 4\end{matrix}\right]-3\left[\begin{matrix}2 & 0 \\ 3 & -4\end{matrix}\right]\] \[-1\left[\begin{matrix}0 & 1 \\ -4 & 4\end{matrix}\right] =-1*[0*4-(-4)*1]=-4\] \[-2\left[\begin{matrix}2 & 1 \\ 3 & 4\end{matrix}\right]=-2*[2*4-3]=-10\] \[-3\left[\begin{matrix}2 & 0 \\ 3 & -4\end{matrix}\right]=-3[2*(-4)-3*0]=24\] -4 - 10 +24=10

OpenStudy (amistre64):

-(0 +4 +16) -1 2 -3 -1 2 2 0 1 2 0 3 -4 4 3 -4 +(0 +6 +24)

OpenStudy (amistre64):

30-20=10 yep

OpenStudy (amistre64):

surdawi has the usual method that is taught ....

OpenStudy (anonymous):



OpenStudy (anonymous):

just like amistre64 do